# Dual metrics on the boundary of strictly polyhedral hyperbolic 3-manifolds

### v4

@@ 2.5 Topology on the space of strictly polyhedral metrics
For $\overline{g}\in\mathcal{CH}(N)$ consider an oriented plane in $\widetilde{N}(\overline{g})=\mathbb{H}^{3}$. It gives rise to an immersed oriented plane in $N(\overline{g})$. We say that an oriented plane is positive if it is at positive distance from $\widetilde{C}(\overline{g})$ and is oriented outwards $\widetilde{C}(\overline{g})$. We say the same about the immersed image of the plane. We denote by $\mathcal{CH}(N,V)$ the space of convex cocompact hyperbolic metrics on $N$ equipped with a set of immersed positive planes marked by $V$, considered up to isometry sending marked planes to marked planes, preserving the marking and isotopic to the identity. When we say that the planes are marked by $V$, we mean that the planes in the geometric end of $N(\overline{g})$ corresponding to $S_{j}$ are marked by points from $S_{j}$. The elements of $\mathcal{CH}(N,V)$ are represented by pairs $(\overline{g},f)$, where $\overline{g}$ is a convex cocompact hyperbolic metric on $N$ and $f$ is a bijection from $V$ onto a set of immersed positive planes. We will write $f_{v}$ instead of $f(v)$. We are going now to endow the space $\mathcal{CH}(N,V)$ with a natural topology.
@@
For $g\in\mathcal{P}(M,V)$ every face gives rise to an immersed plane, which we consider oriented outwards $M(g)$. As shown in Section 2.4, the lift of any such plane does not contain points of $\Lambda(\overline{g})$ at infinity, hence, these planes are at positive distance from $C(\overline{g})$. This allows us to consider $\mathcal{P}(M,V)$ as a subset of $\mathcal{CH}(N,V)$. Clearly, it is open when it is non-empty. We endow $\mathcal{P}(M,V)$ with the induced topology.
@@
To every pair $(\overline{g},f)\in\mathcal{CH}(N,V)$, we correspond the intersection $K$ of all the closed negative immersed half-spaces in $N(\overline{g})$ bounded by the immersed marked planes. We denote by $\overline{\mathcal{P}}(M,V)$ the subset of $\mathcal{CH}(N,V)$ such that (1) $K$ is homeomorphic to $M$; (2) $f$ is injective; (3) each marked plane is supporting to $K$. We remark that here we use a non-intuitive notation as $\overline{\mathcal{P}}(M,V)$ is not a topological closure of $\mathcal{P}(M,V)$: the latter also contains configurations with non-injective markings and with cusps. Our motivation for this notation is that an element of $\overline{P}(M,V)$ determines a dual metric in $\overline{\mathcal{D}}_{c}(\partial M,V)$. We denote by $g$ the metric on $M$ induced by a homeomorphism between $M$ and $K$, where a homeomorphism is chosen so that when it is composed with our fixed homeomorphism $N\cong{\rm int}(M)$, we get the identity on the fundamental groups. The metric $g$ is determined up to isotopy. We will represent the elements of $\overline{\mathcal{P}}(M,V)$ by pairs $(g,f)$. For every such pair there is a subset $V(g)\subseteq V$ where $v\in V(g)$ if and only if the intersection of the plane $f_{v}$ with $\partial M(g)$ has non-empty relative interior, i.e., corresponds to a face. A pair $(g,f)$ also determines a class in $\mathcal{P}(M,V(g))$.
@@
Pick a basis for $\pi_{1}(M)$, let $m$ be its size. Consider three vertices of $\widetilde{M}(g)^{*}$ and act on them by each basis element of $\pi_{1}(M)$. Additionally, for each $v\in V$ take a vertex $\widetilde{v}$ of $\widetilde{M}(g)^{*}$ corresponding to $v$. Let $\widetilde{V}_{0}\subset d\mathbb{S}^{3}$ be the set of these representatives together with all the images of the vertices above, so it is a subset of vertices of $\widetilde{M}(g)^{*}$ of size $n+3m$. Since each element of $G$ is determined by its action on three non-collinear points, one can see that the position of the set $\widetilde{V}_{0}$ in $d\mathbb{S}^{3}$ completely determines the representation $\rho_{\overline{g}}$ and the set $\widetilde{M}(g)^{*}$. Thus, it determines $g$. Moreover, if we vary the position of $\widetilde{V}_{0}$ small enough, we can reconstruct a new discrete and faithful representation $\rho^{\prime}$ of $\pi_{1}(M)$, and all points of $\widetilde{V}_{0}$ will still be in convex position with respect to their orbits under $\rho^{\prime}$. This produces a metric $g^{\prime}\in\mathcal{P}(M,V)$ and all sufficiently close metrics to $g$ arise like this. This shows that there is a neighborhood $Y$ of $g$ in $\mathcal{P}(M,V)$, a neighborhood $X\subset(d\mathbb{S}^{3})^{\widetilde{V}_{0}}$ of the positions of $\widetilde{V}_{0}$ and a smooth submersion $X\rightarrow Y$. It is sufficient to show that the edge-lengths of $\mathcal{T}$ are $C^{1}$-functions in $X$.
@@
Let $Q\subset d\mathbb{S}^{3}$ be a convex geodesic space-like polygon with more than three vertices, $V_{Q}$ be the set of its vertices, $e$ be its interior edge with endpoints $v_{1}$ and $v_{2}$. We denote by $s$ the de Sitter distance between $v_{1}$ and $v_{2}$. When we slightly vary the positions of $V_{Q}$ and take the convex hull, we may obtain a 3-dimensional polytope, whose boundary is naturally divided into two parts, both of which can be considered as variations of $Q$. Let us restrict the attention to one of the parts, we call it upper. If the deformation is small enough, this polytope has only space-like faces, and the total angles at the vertices of the upper part in the intrinsic metric stay less than $\pi$. Let $X_{Q}\subset(d\mathbb{S}^{3})^{V_{Q}}$ be a small enough neighborhood of the initial position of $V_{Q}$, for which this holds. Then we can identify a deformation of $e$ in the upper part, which is a broken line if $e$ is not an edge of the obtained polytope. Consider the length $l$ of $e$ as a function over $X_{Q}$. We need to show that it is a $C^{1}$-function.
@@
Now let $\mathcal{T}_{1},\ldots,\mathcal{T}_{r}$ be all possible triangulations of $Q$. Fix some $\mathcal{T}_{i}$ and consider deformations of $Q$ triangulated with $\mathcal{T}_{i}$ under small variations of $V_{Q}$. Here the obtained polyhedral surface may become non-convex. We assume that $X_{Q}$ is small enough that for all $\mathcal{T}_{i}$ all faces of the deformed surface are space-like and all angles of the vertices in the intrinsic metric are less than $\pi$. Let $l_{i}$ be the length of $e$ under such deformations. Clearly, it is a smooth function in $X_{Q}$ as all the edge-lengths of $\mathcal{T}_{i}$ are smooth functions. For each $v\in V_{Q}$ we choose coordinates such that two of them, called horizontal, are the coordinates of the projection of $v$ to the plane containing $Q$ in the initial moment, and the third, called vertical, is the oriented length of a perpendicular from $v$ to this plane. The latter is a purely imaginary number, but we consider it as a real.
@@
Thus, for fixed $c\neq 0$, $\partial b/\partial a=0$ at $a=0$. Thus, the vertical partial derivatives of all edge-lengths of $\mathcal{T}_{i}$ are zero. In turn, this shows that the vertical partial derivatives of $l_{i}$ are zero too. Now we remark that all horizontal partial derivatives of $l_{i}$ in the initial moment coincide with the horizontal partial derivatives of $s$, the de Sitter distance between the endpoints of $e$.
We return to the convex hull of $V_{Q}$. The domain $X_{Q}$ admits a piecewise-smooth decomposition into cells $E_{1},\ldots,E_{r}$ corresponding to the face triangulations of the upper part of ${\rm conv}(V_{Q})$. The function $l$ coincides with $l_{i}$ over $E_{i}$. The argument above shows that in the initial moment the differential of each $l_{i}$ coincides with the differential of $s$. Thus, $l$ is differentiable in the initial moment. Moreover, it follows that at any point of $X_{Q}$ belonging to several $E_{i}$, the differentials of the respective $l_{i}$ coincide. This implies that $l$ is differentiable over $X_{Q}$. Since all partial derivatives of $l_{i}$ are continuous over $E_{i}$ and coincide with the partial derivatives of $E_{j}$ at the common points of $E_{i}$ and $E_{j}$, we see that $l$ is $C^{1}$. This finishes the proof. ∎
@@ 3 Infinitesimal rigidity
Take two copies of $M$ and identify the corresponding boundary points reversing the orientation. We obtain a closed 3-manifold $M^{D}$ called the double of $M$. A strictly polyhedral hyperbolic metric $g$ on $M$ produces a metric $g^{D}$ on $M^{D}$, which is a hyperbolic cone-metric with the singular locus coming from the edges of $M(g)$ and with all cone-angles smaller than $2\pi$. We refer to [30], [70] or [40] for introductions to hyperbolic cone-3-manifolds. To uniformize the notation, we will write $M^{D}(g^{D})$ for the hyperbolic cone-3-manifold $(M^{D},g^{D})$. We denote by $\Sigma(g^{D})$ the singular locus of $g^{D}$ and denote by $L(g^{D}):=M^{D}(g^{D})\backslash\Sigma(g^{D})$ the smooth part of $M^{D}(g^{D})$ equipped with the induced metric.
@@
Recall that the group $G$ of the orientation-preserving isometries of $\mathbb{H}^{3}$ is isomorphic to $PSL(2,\mathbb{C})$. It is convenient now to consider its universal cover $\widetilde{G}\cong SL(2,\mathbb{C})$. As usual, a holonomy map $\pi_{1}(L)\rightarrow G$, defined up to conjugation, is assigned to every hyperbolic metric from $\mathcal{H}(L)$. It can be lifted to a holonomy map $\pi_{1}(L)\rightarrow\widetilde{G}$ by a result of Culler [16]. Again, we consider the space of representations $\mathcal{R}(\pi_{1}(L),\widetilde{G})$ and its quotient $\mathcal{X}(\pi_{1}(L),\widetilde{G})$ by conjugation of $\widetilde{G}$. The characters of holonomies determine a map $\mathcal{H}(L)\rightarrow\mathcal{X}(\pi_{1}(L),\widetilde{G})$. A version of the Ehresmann-Thurston theorem [10, Theorem 1.7.1] (see also [15, Section 5.3]) shows that this map is a local homeomorphism. The works of Montcouquiol and Weiss [44, 71] show that around the characters of cone-manifolds, $\mathcal{X}(\pi_{1}(L),\widetilde{G})$ locally has the topology of a complex manifold of complex dimension $3(n-k)$ (recall that $k$ is the Euler characteristic of $\partial M$). We endow $\mathcal{H}(L)$ with the pull-back of this smooth structure.
@@
Now suppose that there is $x\in T_{g}\mathcal{P}(M,V)$ such that $dD(x)=0$. Denote by $A$ the closure of $\iota({\rm int}(M))$ in $L$. The inclusion $A\hookrightarrow L$ is a retract. Indeed, the complement $L\backslash A$ is naturally homeomorphic to ${\rm int}(M)$. We send every point of $L\backslash A$ to the corresponding point of $\iota({\rm int}(M))\subset A$, this produces a retraction map. Hence, it induces an injection of fundamental groups $\pi_{1}(M)\cong\pi_{1}(A)\hookrightarrow\pi_{1}(L)$. Thus, $x$ induces zero variation at the first order on the holonomy representation $\rho_{\overline{g}}$ of $\pi_{1}(M)$, and consequently, induces zero variation at the first order on $\overline{g}$.
@@
where $H^{1}(\pi_{1}(L),{\rm Ad}\circ\rho)$ is the first cohomology group of $\pi_{1}(L)$ with respect to the representation ${\rm Ad}\circ\rho$ on the Lie algebra $\mathfrak{g}=\mathfrak{sl}(2,\mathbb{C})$ of $\widetilde{G}$, and ${\rm Ad}$ is the adjoint representation of $\widetilde{G}$ on $\mathfrak{g}$. See the details in [53, Chapter VI][30, Section 4][70, Section 6.2].
Let $\widetilde{E}\rightarrow\widetilde{L}$ be the trivial bundle, where $\widetilde{E}$ is identified with $\widetilde{L}\times\mathfrak{g}$ and $\widetilde{L}$ is endowed with the metric $g^{D}$. The trivialization furnishes the bundle with a flat connection. It is also endowed with a natural, yet non-parallel metric, see, e.g., [30, Section 1]. The group $\pi_{1}(M)$ acts on $\widetilde{E}$ via the representation ${\rm Ad}\circ\rho$. By taking the quotient, we obtain a flat bundle $E\rightarrow L$ of the infinitesimal isometries of $L$. The flat connection induces an exterior derivative $d^{E}$ on the complex $\Omega^{\bullet}(L,E)$ of $E$-valued differential forms on $L$. Consider the respective cohomology complex $H^{\bullet}(L,E)$. An analogue of the de Rham Theorem in this setting establishes that
@@
The details can be found in [53, Chapter VII][30, Section 1][70, Section 6]. We denote by $H^{1}_{L^{2}}(L,E)$ the first $L^{2}$-cohomologies, i.e., the cohomologies of the subcomplex of $L^{2}$-bounded differential forms with $L^{2}$-bounded differential.
@@
Now let $\Xi$ be the smooth part of an oriented convex spherical cone-surface with at least 3 vertices. For $\alpha_{1},\alpha_{2},\alpha_{3}\in(0,2\pi)$ we denote by $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$ the smooth part of a (unique) spherical cone-surface with exactly three cone-angles of given values, if it exists. (It is easy to see that it exists if and only if the sum of the two largest angles is smaller than the smallest one plus $\pi$.) Let $\mathcal{P}$ be a pair of pants decomposition of $\Xi$. Assume that the holonomy angle of every curve in $\mathcal{P}$ is $<2\pi$ and for every pair of pants $\Xi^{\prime}$ of $\mathcal{P}$ there exists the respective surface $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$, where the angles are either the holonomy angles of the boundary curves of $\Xi^{\prime}$ or the cone-angles of vertices. A developing map allows to immerse the surface $\Xi^{\prime}$ into $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$. Assume additionally that this immersion is an embedding for every pair of pants $\Xi^{\prime}$. If all our assumptions hold, we say that $\mathcal{P}$ is suitable. Now we can formulate our main technical lemma
@@
We think that the condition that $\mathcal{P}_{u}$ are suitable is redundant. The suitability, however, allows to use directly subtle analytic bounds from [71], namely, Lemma 3.9 below. Moreover, in our situation it is straightforward that the pair of pants decompositions under consideration are suitable. Without suitability we may try to develop a small neighbourhood of each curve $\gamma$ to a respective spherical suspension and perform there a similar analysis as it is done below. This should allow us to obtain slightly worse (due to spectral properties of spherical suspensions, see [71, Lemma 3.18]), but still sufficient bounds. We choose not to pursue this path.
@@
Denote $d(\chi\circ D)(x)\in H^{1}(L,E)$ by $y$. As $y$ is tangent to the $D$-image of $\mathcal{P}(M,V)$, in particular, it is tangent to $\mathcal{C}(M^{D})$. Then for all meridians $\gamma_{e}$ we get ${\rm Im}(d{\rm tr}_{\gamma_{e}})(y)=0$. Next, as $d\mathcal{I}_{V}(x)=0$, then for the dihedral angle $\theta_{e}$ of each edge we have $d\theta_{e}(x)=0$. Thus, ${\rm Re}(d{\rm tr}_{\gamma_{e}})(y)=0$, so $d{\rm tr}_{\gamma_{e}}(y)=0$.
Consider a vertex $u$ of $\Sigma$. Look at $u$ also as at a vertex of $M$. Its spherical link in $M$ is a convex spherical polygon $Z$, and its spherical link in $L$ is a convex spherical cone-surface $Z^{D}$, which is the double of $Z$. Denote also by $Z^{*}$ the dual spherical polygon to $Z$. The tangent vector $x$ induces the decomposition of $Z^{*}$ into convex polygons (without adding new vertices), which we extend to a triangulation. For each interior edge of this triangulation we consider an arbitrary smooth curve in $Z$ connecting the respective edges of $Z$ and orthogonal to them, such that these curves are disjoint for different interior edges of the triangulation. This produces a decomposition of $Z$ into curvilinear polygons. The double of this decomposition is a pair of pants decomposition of the smooth part of $Z^{D}$, which projects to a pair of pants decomposition $\mathcal{P}_{u}$ of $\Xi_{u}$. Note that it is suitable: this is easy to see since every curvilinear polygon $C$ in $Z$ belongs to the convex spherical triangle formed by the extensions of those edges of $C$ that come from the edges of $Z$ (note that there are always three such edges). Since $d\mathcal{I}_{V}(x)=0$, by construction, $d{\rm tr}_{\gamma}(y)=0$ for every $\gamma\in\mathcal{P}_{u}$.
@@
Our proof of Lemma 3.4 will employ the Bochner Technique and will follow the proof of [71, Theorem 3.15], though there will be some differences. We need to make more preparations. Pick some $\Xi_{u}=\Xi$ with the restriction of the bundle $E$ to it and a $\rho$-admissible pair of pants decomposition $\mathcal{P}_{u}=\mathcal{P}$. Let $\mathcal{Q}_{u}=\mathcal{Q}$ be a set consisting of one peripheral curve in $\Xi$ for each cone-point. From [71, Lemma 3.10]
@@
Here and in what currently follows the kernels are taken in $H^{1}(\Xi,E)$. Due to [71, Lemma 3.9], we have
@@
By Lemma 3.3, $\{d{\rm tr}_{\gamma}\}_{\gamma\in\mathcal{Q}}$ are $\mathbb{C}$-linearly independent in $H^{1}(\Xi,E)^{*}$, where *** means the dual vector space. This, (3.1) and (3.2) show that

### v5

@@ 2.5 Topology on the space of strictly polyhedral metrics
For $\overline{g}\in\mathcal{CH}(N)$ consider an oriented plane in $\widetilde{N}(\overline{g})=\mathbb{H}^{3}$. It gives rise to an immersed oriented plane in $N(\overline{g})$. We say that an oriented plane is positive if it is at positive distance from $\widetilde{C}(\overline{g})$ and is oriented outwards $\widetilde{C}(\overline{g})$. We say the same about the immersed image of the plane. We denote by $\mathcal{CH}(N,V)$ the space of convex cocompact hyperbolic metrics on $N$ equipped with a set of immersed positive planes marked by $V$, considered up to isometries sending marked planes to marked planes, preserving the marking and isotopic to the identity. When we say that the planes are marked by $V$, we mean that the planes in the geometric end of $N(\overline{g})$ corresponding to $S_{j}$ are marked by points from $S_{j}$. The elements of $\mathcal{CH}(N,V)$ are represented by pairs $(\overline{g},f)$, where $\overline{g}$ is a convex cocompact hyperbolic metric on $N$ and $f$ is a map from $V$ onto a set of immersed positive planes. We will write $f_{v}$ instead of $f(v)$. We are going now to endow the space $\mathcal{CH}(N,V)$ with a natural topology.
@@
For $g\in\mathcal{P}(M,V)$ every face gives rise to an immersed plane, which we consider oriented outwards $M(g)$. As shown in Section 2.4, the lift of any such plane does not contain points of $\Lambda(\overline{g})$ at infinity, hence, these planes are at positive distance from $C(\overline{g})$. This allows us to consider $\mathcal{P}(M,V)$ as a subset of $\mathcal{CH}(N,V)$. Clearly, it is open. We endow $\mathcal{P}(M,V)$ with the induced topology.
@@
For every pair $(\overline{g},f)\in\mathcal{CH}(N,V)$ we consider the intersection $K$ of all the closed negative immersed half-spaces in $N(\overline{g})$ bounded by the immersed marked planes. We denote by $\overline{\mathcal{P}}(M,V)$ the subset of $\mathcal{CH}(N,V)$ such that (1) $K$ is homeomorphic to $M$; (2) $f$ is injective; (3) each marked plane is supporting to $K$. We remark that here we use a non-intuitive notation as $\overline{\mathcal{P}}(M,V)$ is not a topological closure of $\mathcal{P}(M,V)$: the latter also contains configurations with non-injective markings and with cusps. Our motivation for this notation is that an element of $\overline{P}(M,V)$ determines a dual metric in $\overline{\mathcal{D}}_{c}(\partial M,V)$. We denote by $g$ the metric on $M$ induced by a homeomorphism between $M$ and $K$, where a homeomorphism is chosen so that when it is composed with our fixed homeomorphism $N\cong{\rm int}(M)$, we get the identity on the fundamental groups. The metric $g$ is determined up to isotopy. We will represent the elements of $\overline{\mathcal{P}}(M,V)$ by pairs $(g,f)$. For every such pair there is a subset $V(g)\subseteq V$ where $v\in V(g)$ if and only if the intersection of the plane $f_{v}$ with $\partial M(g)$ has non-empty relative interior, i.e., corresponds to a face. A pair $(g,f)$ also determines a class in $\mathcal{P}(M,V(g))$.
@@
Let $Q\subset d\mathbb{S}^{3}$ be a convex geodesic space-like polygon with more than three vertices representing a face of $\widetilde{M}(g)$, $V_{Q}$ be the set of its vertices, $\widetilde{e}$ be its interior edge with endpoints $v_{1}$ and $v_{2}$ representing an edge $e$ of $\mathcal{T}$. We denote by $s$ the de Sitter distance between $v_{1}$ and $v_{2}$. When we slightly vary the positions of $V_{Q}$ and take the convex hull, we may obtain a 3-dimensional polytope, whose boundary is naturally divided into two parts, both of which can be considered as variations of $Q$. Let us restrict the attention to one of the parts, we call it upper. If the deformation is small enough, this polytope has only space-like faces, and the total angles at the vertices of the upper part in the intrinsic metric stay less than $\pi$. Let $X_{Q}\subset(d\mathbb{S}^{3})^{V_{Q}}$ be a small enough neighborhood of the initial position of $V_{Q}$, for which this holds. Then we can identify a deformation of $\widetilde{e}$ in the upper part, which is a broken line if $\widetilde{e}$ is not an edge of the obtained polytope. Consider the length $\widetilde{l}$ of $\widetilde{e}$ as a function over $X_{Q}$. There exists a small neighborhood $Y$ of $g$ in $\mathcal{P}(M,V)$, over which the length $l$ of $e$ can be defined, and a smooth map $q:Y\rightarrow X_{Q}$ such that $l=\widetilde{l}\circ q$. Thus, it is enough to show that $\widetilde{l}$ is a $C^{1}$-function.
@@
Now let $\mathcal{T}_{1},\ldots,\mathcal{T}_{r}$ be all possible triangulations of $Q$. Fix some $\mathcal{T}_{i}$ and consider deformations of $Q$ triangulated with $\mathcal{T}_{i}$ under small variations of $V_{Q}$. Here the obtained polyhedral surface may become non-convex. We assume that $X_{Q}$ is small enough that for all $\mathcal{T}_{i}$ all faces of the deformed surface are space-like and all angles of the vertices in the intrinsic metric are less than $\pi$. Let $l_{i}$ be the length of $\widetilde{e}$ under such deformations. Clearly, it is a smooth function in $X_{Q}$ as all the edge-lengths of $\mathcal{T}_{i}$ are smooth functions. For each $v\in V_{Q}$ we choose coordinates such that two of them, called horizontal, are the coordinates of the projection of $v$ to the plane containing $Q$ in the initial moment, and the third, called vertical, is the oriented length of a perpendicular from $v$ to this plane. Note that since the plane is space-like, the perpendicular is time-like, but here we consider its length as a real number (with the sign depending on the position of $v$ with respect to the plane).
@@
Thus, for fixed $c\neq 0$, $\partial b/\partial a=0$ at $a=0$. Thus, the vertical partial derivatives of all edge-lengths of $\mathcal{T}_{i}$ are zero. In turn, this shows that the vertical partial derivatives of $l_{i}$ are zero too. Now we remark that all horizontal partial derivatives of $l_{i}$ in the initial moment coincide with the horizontal partial derivatives of $s$, the de Sitter distance between the endpoints of $\widetilde{e}$.
We return to the convex hull of $V_{Q}$. The domain $X_{Q}$ admits a piecewise-smooth decomposition into cells $E_{1},\ldots,E_{r}$ corresponding to the face triangulations of the upper part of ${\rm conv}(V_{Q})$. The function $\widetilde{l}$ coincides with $l_{i}$ over $E_{i}$. The argument above shows that in the initial moment the differential of each $l_{i}$ coincides with the differential of $s$. Thus, $\widetilde{l}$ is differentiable in the initial moment. Moreover, it follows that at any point of $X_{Q}$ belonging to several $E_{i}$, the differentials of the respective $l_{i}$ coincide. This implies that $\widetilde{l}$ is differentiable over $X_{Q}$. Since all partial derivatives of $l_{i}$ are continuous over $E_{i}$ and coincide with the partial derivatives of $E_{j}$ at the common points of $E_{i}$ and $E_{j}$, we see that $\widetilde{l}$ is $C^{1}$. This finishes the proof. ∎
@@ 3 Infinitesimal rigidity
Take two copies of $M$ and identify the corresponding boundary points reversing the orientation. We obtain a closed 3-manifold $M^{D}$ called the double of $M$. A strictly polyhedral hyperbolic metric $g$ on $M$ produces a metric $g^{D}$ on $M^{D}$, which is a hyperbolic cone-metric with the singular locus coming from the edges of $M(g)$ and with all cone-angles smaller than $2\pi$. We refer to [32], [73] or [42] for introductions to hyperbolic cone-3-manifolds. To uniformize the notation, we will write $M^{D}(g^{D})$ for the hyperbolic cone-3-manifold $(M^{D},g^{D})$. We denote by $\Sigma(g^{D})$ the singular locus of $g^{D}$ and denote by $L(g^{D}):=M^{D}(g^{D})\backslash\Sigma(g^{D})$, the smooth part of $M^{D}(g^{D})$ equipped with the induced metric.
@@
Recall that the group $G$ of the orientation-preserving isometries of $\mathbb{H}^{3}$ is isomorphic to ${\rm PSL}(2,\mathbb{C})$. It is convenient now to consider its universal cover $\widetilde{G}\cong{\rm SL}(2,\mathbb{C})$. As usual, a holonomy map $\pi_{1}(L)\rightarrow G$, defined up to conjugation, is assigned to every hyperbolic metric from $\mathcal{H}(L)$. It can be lifted to a holonomy map $\pi_{1}(L)\rightarrow\widetilde{G}$ by a result of Culler [16]. Again, we consider the space of representations $\mathcal{R}(\pi_{1}(L),\widetilde{G})$ and its quotient $\mathcal{X}(\pi_{1}(L),\widetilde{G})$ by conjugation of $\widetilde{G}$. The characters of holonomies determine a map $\mathcal{H}(L)\rightarrow\mathcal{X}(\pi_{1}(L),\widetilde{G})$. A version of the Ehresmann-Thurston theorem [10, Theorem 1.7.1] (see also [15, Section 5.3]) shows that this map is a local homeomorphism. The works of Montcouquiol and Weiss [46, 74] show that around the characters of cone-manifolds, $\mathcal{X}(\pi_{1}(L),\widetilde{G})$ locally has the topology of a complex manifold of complex dimension $3(n-k)$ (recall that $k$ is the Euler characteristic of $\partial M$). We endow $\mathcal{H}(L)$ with the pull-back of this smooth structure.
@@
Now suppose that there is $x\in T_{g}\mathcal{P}(M,V)$ such that $dD(x)=0$. Denote by $A$ the closure of $\iota({\rm int}(M))$ in $L$. The inclusion $A\hookrightarrow L$ is a retract. Indeed, the complement $L\backslash A$ is naturally homeomorphic to ${\rm int}(M)$. We send every point of $L\backslash A$ to the corresponding point of $\iota({\rm int}(M))\subset A$, this produces a retraction map. Hence, it induces an injection of fundamental groups $\pi_{1}(M)\cong\pi_{1}(A)\hookrightarrow\pi_{1}(L)$. Thus, $x$ induces zero variation at the first order on the holonomy representation $\rho_{\overline{g}}$ of $\pi_{1}(M)$, and, consequently, induces zero variation at the first order on $\overline{g}$.
@@
where $H^{1}(\pi_{1}(L),{\rm Ad}\circ\rho)$ is the first cohomology group of $\pi_{1}(L)$ with respect to the representation ${\rm Ad}\circ\rho$ on the Lie algebra $\mathfrak{g}=\mathfrak{sl}(2,\mathbb{C})$ of $\widetilde{G}$, and ${\rm Ad}$ is the adjoint representation of $\widetilde{G}$ on $\mathfrak{g}$. See the details in [55, Chapter VI][32, Section 4][73, Section 6.2].
Let $\widetilde{E}\rightarrow\widetilde{L}$ be the trivial bundle, where $\widetilde{E}$ is identified with $\widetilde{L}\times\mathfrak{g}$ and $\widetilde{L}$ is endowed with the metric $g^{D}$. The trivialization furnishes the bundle with a flat connection. It is also endowed with a natural, yet non-parallel metric, see, e.g., [32, Section 1]. The group $\pi_{1}(M)$ acts on $\widetilde{E}$ via the representation ${\rm Ad}\circ\rho$. By taking the quotient, we obtain a flat bundle $E\rightarrow L$ of the infinitesimal isometries of $L$. The flat connection induces an exterior derivative $d^{E}$ on the complex $\Omega^{\bullet}(L,E)$ of $E$-valued differential forms on $L$. Consider the respective cohomology complex $H^{\bullet}(L,E)$. An analogue of the de Rham Theorem in this setting establishes that
@@
The details can be found in [55, Chapter VII][32, Section 1][73, Section 6]. We denote by $H^{1}_{L^{2}}(L,E)$ the first $L^{2}$-cohomologies, i.e., the cohomologies of the subcomplex of $L^{2}$-bounded differential forms with $L^{2}$-bounded differential.
@@
For $\alpha_{1},\alpha_{2},\alpha_{3}\in(0,2\pi)$ we denote by $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$ the smooth part of a (unique) spherical cone-surface with exactly three cone-angles of given values, if it exists. (It is easy to see that it exists if and only if the sum of the two largest angles is smaller than the smallest one plus $\pi$.) Now let $\Xi$ be the smooth part of an oriented convex spherical cone-surface with at least 3 vertices. Let $\mathcal{P}$ be a pair of pants decomposition of $\Xi$. Assume that the holonomy angle of every curve in $\mathcal{P}$ is $<2\pi$ and for every pair of pants $\Xi^{\prime}$ of $\mathcal{P}$ there exists the respective surface $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$, where the angles are either the holonomy angles of the boundary curves of $\Xi^{\prime}$ or the cone-angles of vertices. A developing map allows to immerse the surface $\Xi^{\prime}$ into $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$. Assume additionally that this immersion is an embedding for every pair of pants $\Xi^{\prime}$. If all our assumptions hold, we say that $\mathcal{P}$ is suitable. Now we can formulate our main technical lemma
@@
We think that the condition that $\mathcal{P}_{u}$ are suitable is redundant. The suitability, however, allows to use directly subtle analytic bounds from [74], namely, Lemma 3.9 below. Moreover, in our situation it is straightforward that the pair of pants decompositions under consideration are suitable. Without suitability we may try to develop a small neighbourhood of each curve $\gamma$ to a respective spherical suspension and perform there a similar analysis as it is done below. This should allow us to obtain slightly worse (due to spectral properties of spherical suspensions, see [74, Lemma 3.18]), but still sufficient bounds. We choose not to pursue this path.
@@
Denote $dD(x)$ by $y$, considered as an element of $H^{1}(L,E)$. As $y$ is tangent to the $D$-image of $\mathcal{P}(M,V)$, in particular, it is tangent to $\mathcal{C}(M^{D})$. Then for all meridians $\gamma_{e}$ we get ${\rm Im}(d{\rm tr}_{\gamma_{e}})(y)=0$. Next, as $d\mathcal{I}_{V}(x)=0$, then for the dihedral angle $\theta_{e}$ of each edge we have $d\theta_{e}(x)=0$. Thus, ${\rm Re}(d{\rm tr}_{\gamma_{e}})(y)=0$, so $d{\rm tr}_{\gamma_{e}}(y)=0$.
Consider a vertex $u$ of $\Sigma$. Look at $u$ also as at a vertex of $M$. Its spherical link in $M$ is a convex spherical polygon $Z$, and its spherical link in $M^{D}$ is a convex spherical cone-surface $Z^{D}$, which is the double of $Z$. Denote also by $Z^{*}$ the dual spherical polygon to $Z$. The tangent vector $x$ induces the decomposition of $Z^{*}$ into convex polygons (without adding new vertices), which we extend to a triangulation. For each interior edge of this triangulation we consider an arbitrary smooth curve in $Z$ connecting the respective edges of $Z$ and orthogonal to them, such that these curves are disjoint for different interior edges of the triangulation. This produces a decomposition of $Z$ into curvilinear polygons. The double of this decomposition is a pair of pants decomposition of the smooth part of $Z^{D}$, which gives a pair of pants decomposition $\mathcal{P}_{u}$ of $\Xi_{u}$. Note that it is suitable: this is easy to see since every curvilinear polygon $C$ in $Z$ belongs to the convex spherical triangle formed by the extensions of those edges of $C$ that come from the edges of $Z$ (note that there are always three such edges). Since $d\mathcal{I}_{V}(x)=0$, by construction, $d{\rm tr}_{\gamma}(y)=0$ for every $\gamma\in\mathcal{P}_{u}$.
@@
Our proof of Lemma 3.4 will employ the Bochner Technique and will follow the proof of [74, Theorem 3.15], though there will be some differences. We need to make more preparations. Pick some $\Xi_{u}=\Xi$ with the restriction of the bundle $E$ to it and a $\rho$-admissible pair of pants decomposition $\mathcal{P}_{u}=\mathcal{P}$. Let $\mathcal{Q}_{u}=\mathcal{Q}$ be a set consisting of one peripheral curve in $\Xi$ for each cone-point. From [74, Lemma 3.10]
@@
Here and in what currently follows the kernels are taken in $H^{1}(\Xi,E)$. Due to [74, Lemma 3.9], we have
@@
By Lemma 3.3, $\{d{\rm tr}_{\gamma}\}_{\gamma\in\mathcal{Q}}$ are $\mathbb{C}$-linearly independent in $H^{1}(\Xi,E)^{*}$, where * means the dual vector space. This, (3.1) and (3.2) show that